OpenStudy (anonymous):

Write an equation in point-slope form of the line through point J(4, –4) with slope 4.

OpenStudy (ipwnbunnies):

\[Y-Y1 = m(X-X1)\]

OpenStudy (ipwnbunnies):

That's point-slope form, where (X1.Y1) is the given point.

OpenStudy (anonymous):

so y-4=4x+4 ?

OpenStudy (ipwnbunnies):

Err no try again. Also, 'm' is the slope of the line. Also don't forget youur Y value is actually negative.

OpenStudy (anonymous):

or y+4=4x-4

OpenStudy (ipwnbunnies):

Eh, almost. You got the left part right. This is what the right should look like: 4(x-4) You have to distribute the 4 to each component.

OpenStudy (anonymous):

-4(x-4) ?

OpenStudy (ipwnbunnies):

Err, no lol. We're keeping 4(x-4) But you have to distribute the 4 to each term. 4(x-4) = 4*x - 4*4 = 4x - 16. y + 4 = 4x - 16

OpenStudy (anonymous):

well that's not any of my choices

OpenStudy (ipwnbunnies):

Ehh, what are your choices? o.o

OpenStudy (ipwnbunnies):

You have to reduce it to slope-intercept form, perhaps? Which means getting y by itself?

OpenStudy (anonymous):

a. y+4=4(x+4) b. y+4=4(x-4) c. y-4=4(x+4) d. y+4=-4(x-4)

OpenStudy (ipwnbunnies):

Oh yeah, it's B. y + 4 = 4(x-4)

OpenStudy (ipwnbunnies):

Das what I showed you. :3

OpenStudy (anonymous):

@Jack1

OpenStudy (ipwnbunnies):

Um...ok.

OpenStudy (anonymous):

i'm stuck between b and c can you help

OpenStudy (anonymous):

that was for jack btw

OpenStudy (ipwnbunnies):

I said it was B! Point (4,-4) , slope of 4 (y-y1) = m(x-x1) (y+4) = 4(x-4)

OpenStudy (jack1):

hey shay, bun's helpin u out here, bear with him

OpenStudy (anonymous):

but you were confusing me taking me all around the world at first

OpenStudy (ipwnbunnies):

Sorry. I was just showing you how to do it.

OpenStudy (anonymous):

when i asked if that was the correct answer way before but okay thanks for your time guys :) much appreciated

OpenStudy (ipwnbunnies):

Ohhh, you didn't put parentheses around your original answer. That's why :o

OpenStudy (anonymous):

ok